# Crowd-FM: Learned Optimal Selection of Conditional Flow Matching-generated Trajectories for Crowd Navigation

### v1

@@ I INTRODUCTION
Mobile robots are still far from achieving effective trajectory modeling that enables them to operate in scenarios characterized by extreme dynamic complexity. Such scenarios are difficult to navigate even for humans, and in these situations, simply prioritizing safety above all else becomes the norm. Classical planners, with their mathematically rigorous models, are effective in ensuring safety but often fail in environments with high uncertainty and variability, as seen in velocity-obstacle–based approaches such as RVO [2]. In contrast, recent approaches based on generative modeling, such as VQ-VAEs [3, 4] or Reinforcement Learning [5, 6, 7], attempt to capture complex behaviors but suffer from limitations in scalability, sample efficiency, and generalization. Furthermore, much of the existing literature demonstrates success only in relatively simple, low-interaction settings [8], while performance degrades substantially in densely interactive environments common in real-world applications.
@@ II RELATED WORKS
Generative Models for Trajectory Planning: Generative modeling has recently emerged as a promising alternative for reinforcement and classical methods. Diffusion-based approaches such as NoMaD [18] learn goal-conditioned distributions for long-horizon planning, while VQ-VAE–based methods like CrowdSurfer [3] and variations with differentiable safety filters [4] have achieved strong results in dense navigation tasks. However, these models suffer from limitations such as computational overhead (diffusion) or due to latent bottlenecks which quantize away detail (VQ-VAE).
Flow-Based Generative Models: Flow-based methods provide an alternative for distribution learning. Lipman et al. [9] introduced Flow Matching as a stable and efficient training paradigm, with later works that improve generalization through minibatch optimal transport [10]. Applications of Conditional Flow Matching have been demonstrated in robotic manipulation from pointclouds [11] and navigation with depth priors [12]. In this work, we extend these ideas to dense crowd navigation, where Flow Matching enables robust trajectory distribution learning, combined with a learned scoring function to boost success.
@@ III-A Trajectory Parameterization
where the matrix $\mathbf{P}$ is a matrix formed by time-dependent polynomial basis functions. The vectors $\mathbf{c}_{x},\mathbf{c}_{y}$ are the coefficients attached to the individual basis functions. We can represent the derivatives like $\dot{x}(t),\ddot{x}(t),\dot{y}(t),\ddot{y}(t)$ in a similar manner as (1) using $\dot{\mathbf{P}}$ and $\ddot{\mathbf{P}}$.
Let $\boldsymbol{\xi}=\begin{bmatrix}\mathbf{c}_{x}\\ \mathbf{c}_{y}\end{bmatrix}$ be the combined representation of the robot’s trajectory.
Our trajectory representation follows previous motion planning works like [3], [4], and [1]. By choosing to generate Bernstein control points rather than a sequence of waypoints, the generative model is not burdened with learning the concepts of smoothness or continuity from data since the mathematical structure of the Bernstein representation guarantees these priors. However, ensuring the resulting trajectories adhere to the robot’s dynamic constraints requires further refinement, which is addressed in III-D2.
Although our choice of Bernstein polynomials is motivated by the same advantages demonstrated in recent diffusion-based works such as GPD [19], the way these representations are used differs substantially. GPD leverages diffusion over Bernstein coefficients with a guided denoising process and a trajectory stitching procedure to ensure feasibility. In contrast, Crowd-FM uses CFM to directly learn the transformation to Bernstein coefficients via ODE integration, enabling faster and more stable inference.
@@ III-B Flow Matching Objective
Following the independent coupling strategy employed in recent works like [12], a conditional probability path $p_{\tau}(\boldsymbol{\xi}|\boldsymbol{\xi_{0}},\boldsymbol{\xi_{1}})$ and a corresponding vector field $u_{\tau}(\boldsymbol{\xi}|\boldsymbol{\xi_{0}},\boldsymbol{\xi_{1}})$ are defined for any pair of samples $\boldsymbol{\xi_{0}}\sim q_{0}(\boldsymbol{\xi})$ and $\boldsymbol{\xi_{1}}\sim q_{1}(\boldsymbol{\xi}).$ This path is a simple linear interpolation in the high-dimensional space of control points:
@@ III-C Flow Matching Model Architecture
This subsection details the Conditional Flow Matching architecture (Fig 2) adapted to learn the complex multi-modal distribution of Bernstein control points that define expert trajectories.
@@ III-C1 Multi-modal Conditioning Encoder
Point Cloud Encoder: For a point cloud input $X_{\mathrm{pcd}}\in\mathbb{R}^{N_{\mathrm{pts}}\times 2}$, where $N_{\mathrm{pts}}$ is the maximum number of point-cloud points, a feature extractor $f_{\mathrm{pcd}}$ composed of a series of 1D convolutions and ReLU activations maps each point to a high-dimensional feature vector. A global max-pooling operation is then applied across all points to yield a permutation-invariant global feature vector, $c_{\mathrm{pcd}}=\max_{i=1...N_{\mathrm{pts}}}f_{\mathrm{pcd}}(X_{\mathrm{pcd}})$. This vector is processed by a multi-layer perceptron (MLP), $g_{\mathrm{pcd}}$, to produce the final static embedding, $C_{\mathrm{static}}=g_{\mathrm{pcd}}(c_{\mathrm{pcd}})$.
Dynamic Obstacle Encoder: The Dynamic Obstacle information is compiled into a specialized tensor for the model to work with. For an input dynamic obstacle tensor $X_{\mathrm{dyn}}\in\mathbb{R}^{N_{\mathrm{obs}}\times 4}$, representing $N_{\mathrm{obs}}$, each with a 4D kinematic state(position and velocity), a dynamic obstacle encoder $f_{\mathrm{dyn}}$ processes them into a high-dimensional feature vector. An MLP converts obstacle features into high-dimensional embeddings, which are augmented with positional embeddings. Subsequently, a minimal Transformer Encoder with 4 heads and w layers employs self-attention to model interactions among the obstacles. Finally, max-pooling aggregates these features into a single vector, which a final MLP processes to produce the dynamic context embedding(D=512)(D=512), $C_{\mathrm{dynamic}}$.
Goal Encoder: The navigation goal is provided as a 2D heading vector, $X_{\mathrm{goal}}\in\mathbb{R}^{2}$. This vector is processed by a dedicated MLP, $g_{\mathrm{goal}}$, to produce a goal embedding: $C_{\mathrm{goal}}=g_{\mathrm{goal}}(\mathrm{X}_{\mathrm{goal}})$.
Multi-modal Fusion: The individual embeddings, $C_{\mathrm{static}}$, $C_{\mathrm{dyn}}$, and $C_{\mathrm{goal}}$, are treated as tokens for a fusion module. The tokens are concatenated to form an input sequence, which is then processed by $F_{\mathrm{fusion}}$, a Transformer Encoder with 8 heads and 3 layers. The Transformer Encoder uses self-attention to produce the final context, a sequence of fused tokens $C=F_{\mathrm{fusion}}(S)$. This sequence captures the complex interplay between the robot’s environment, other moving agents, and its objective. This context is directly used to condition the 1D U-Net as explained below.
@@ III-C2 Conditional U-Net
    The interpolated control point vector $\xi_{\tau}\in\mathbb{R}^{2(n+1)}$.
@@
The U-Net architecture consists of a symmetric encoder-decoder structure with skip connections. The encoder progressively downsamples the 1D representation of the trajectory control points and the decoder upsamples this representation, integrating information from the corresponding encoder layers via skip connections. The flow time $\tau$ is transformed into a sinusoidal time embedding and added to the intermediate features at each resolution level. The context $C$ is injected into the network within the U-Net’s residual blocks, allowing the model to condition the predicted flow on the specific environmental and goal context. The final output of the network is the predicted vector field $v_{\theta}(\boldsymbol{\xi_{\tau}},\tau,C)$, which has the same dimension as the input $\boldsymbol{\xi_{\tau}}$.
@@ III-D1 Guided Integration Process
It is important to note that $p_{k}$ is a function of $\xi_{\tau}$, and is implied that:
$\mathcal{L}_{\text{collision}}=f(\mathbf{p}_{k})\quad\text{and}\quad\mathbf{p}_{k}=g(\xi_{\tau})\implies\mathcal{L}_{\text{collision}}=f(g(\xi_{\tau}))$
@@ III-E2 Training Objective
The Scoring Function is trained using a cross-entropy loss, treating the problem as a K-class classification task. This objective trains the scorer to assign the highest logit to the candidate that best imitates the expert’s path. During inference, it is sufficient to select the trajectory index with the highest score.
@@
A potential concern when framing trajectory selection as a K-class classification problem is the issue of permutation invariance. If the ordering of candidate trajectories were arbitrary and disconnected from their corresponding scores and labels, the cross-entropy loss would yield an inconsistent training signal. However, this concern is addressed by our methodology, wherein the target labels are dynamically generated for each training instance.
@@
Therefore, while the absolute identity of any given index changes between training steps, the relative correspondence between the scores and the dynamically generated target is preserved, ensuring that the cross-entropy loss correctly penalizes the model for failing to assign the highest score to the geometrically optimal trajectory, as seen in Fig 3.
@@ III-F Data Collection
TABLE I: Comparison of navigation success rates across different planners and environments with 35 dynamic agents
Data collection was performed using the Barn environment, to capture challenging static obstacle layout and PEDSIM, to model dense dynamic crowds. The dataset consists of around five hours of navigation data using both the Jackal and Turtlebot2 robot. Two sources of trajectories were included:
@@ IV VALIDATION AND BENCHMARKING
In this section, we detail our extensive evaluation of Crowd-FM. We show qualitative results, conduct quantitative benchmarking and perform ablation studies in challenging simulated environments, comparing our method against three leading baselines: CrowdSurfer[3], DRL-VO[5], and CoHAN2.0[16]. We then validate the practical feasibility of our pipeline in real-world settings.
@@ IV-B Quantitative Analysis
Crowd-FM sets a new standard for performance in crowd navigation, outperforming all baselines in nearly every metric across standard PEDSIM[21] environments (Table I). This particularly improves upon CrowdSurfer, despite their shared use of generative priors. We stress-test both these pipelines in more challenging environments and share our results in Table II.
This improvement stems from two pillars: (1) our CFM model’s near-perfect learning of collision avoidance from data, and (2) exceptional computational speed, enabling full closed-loop inference in <<60ms (CFM alone: $\sim$22ms). This makes the Crowd-FM pipeline both highly effective and an extremely fast solution.
The following subsections are dedicated to ablation studies, especially focusing on the effectiveness of the vanilla CFM-generated trajectories.
@@ IV-B1 Ablation Study without Optimizer Refinement
Here, we test Crowd-FM without the Optimizer refinement step by comparing the trajectories generated by the Vanilla-CFM with another baseline (DRL-VO) and see if the Conditional Flow Matching models the input data distribution well. As is evident from Table III, vanilla CFM alone is able to beat a leading RL-based baseline, DRL-VO, by a fair margin. This shows that CFM alone can model collision avoidance, especially in dense crowds. Qualitative examples of vanilla-CFM trajectories can be found in Fig 5.
@@ IV-B2 Effect of Inference-time Cost Guidance
One of the two ways we employ inference-time refinement of CFM-generated trajectories is the Inference-time Cost Guidance. We test this property of Crowd-FM by disabling Optimizer refinement on the Cumberland and Freiburg environments in Pedsim with 25 agents for 10 trials. Table IV shows a clear improvement in Success rates for the collision-cost guided Flow objective compared to the un-guided one. Additionally, we present qualitative results for the same in Fig 5.
Figure 5: Effect of Inference-time Cost Guidance: Left image shows the trajectories generated by Crowd-FM with Cost Guidance. Right image shows the trajectories generated by Crowd-FM without Cost Guidance. Trajectories generated using collision cost guidance are found to be more controlled near obstacles than the trajectories without it. (Tests done without Optimizer Refinement)
@@ IV-B3 Effect of the Learned Scoring Function
Lower values of HLP indicate that the chosen trajectory is closer to the expert demonstration, and thus more human-like. We compare trajectories selected by our learned Scoring Function against those chosen by predefined, hand-tuned cost functions that primarily optimize for collision avoidance and smoothness. These tests were done in an open-loop setting to take into account the expert human demonstrations. A custom wheelchair prototype was teleoperated in indoor environments amid dense crowds to create the test dataset. We evaluated HLPs for both methods on 10 such scenes and provided our results as a Grouped Bar Chart in Fig 6.

### v2

@@ I INTRODUCTION
Mobile robots are still far from achieving effective trajectory modelling that enables them to operate in scenarios characterized by extreme dynamic complexity. Such scenarios are difficult to navigate even for humans, and in these situations, simply prioritizing safety above all else becomes the norm. Classical planners, with their mathematically rigorous models, are effective in ensuring safety but often fail in environments with high uncertainty and variability, as seen in velocity-obstacle–based approaches such as RVO [19]. In contrast, recent approaches based on generative modelling, such as VQ-VAEs [7, 6] or Reinforcement Learning [20, 2, 12], attempt to capture complex behaviors but suffer from limitations in scalability, sample efficiency, and generalization. Furthermore, much of the existing literature demonstrates success only in relatively simple, low-interaction settings [1], while performance degrades substantially in densely interactive environments common in real-world applications.
@@ II RELATED WORKS
Generative Models for Trajectory Planning: Generative modelling has recently emerged as a promising alternative to reinforcement and classical methods. Diffusion-based approaches such as NoMaD [16] learn goal-conditioned distributions for long-horizon planning, while VQ-VAE–based methods like CrowdSurfer [7] and variations with differentiable safety filters [6] have achieved strong results in dense navigation tasks. However, these models suffer from limitations such as computational overhead (diffusion) or due to latent bottlenecks which quantize away detail (VQ-VAE).
Flow-Based Generative Models: Flow-based methods provide an alternative for distribution learning. Lipman et al. [9] introduced Flow Matching as a stable and efficient training paradigm, with subsequent work improving generalization via minibatch optimal transport [18]. Applications of Conditional Flow Matching have been demonstrated in robotic manipulation from pointclouds [3] and navigation with depth priors [5]. In this work, we extend these ideas to dense crowd navigation, where Flow Matching enables robust trajectory distribution learning, combined with a learned scoring function to boost success.
@@ III-A Trajectory Parameterization
where the matrix $\mathbf{P}$ is a matrix formed by time-dependent polynomial basis functions. The vectors $\mathbf{c}_{x},\mathbf{c}_{y}$ are the coefficients attached to the individual basis functions. We can represent the derivatives like $\dot{x}(t),\ddot{x}(t),\dot{y}(t),\ddot{y}(t)$ in a similar manner as (1) using $\dot{\mathbf{P}}$ and $\ddot{\mathbf{P}}$. The combined trajectory representation is then given as $\boldsymbol{\xi}=\begin{bmatrix}\mathbf{c}_{x}&\mathbf{c}_{y}\end{bmatrix}^{T}$.
Our trajectory representation follows previous motion planning works like [7], [6], and [11]. By choosing to generate Bernstein control points rather than a sequence of waypoints, the generative model is not burdened with learning the concepts of smoothness or continuity from data since the mathematical structure of the Bernstein representation guarantees these priors. However, ensuring the resulting trajectories adhere to the robot’s dynamic constraints requires further refinement, which is addressed in III-D2. Although our choice of Bernstein polynomials is motivated by the same advantages demonstrated in recent diffusion-based works such as GPD [17], the way these representations are used differs substantially. GPD leverages diffusion over Bernstein coefficients with a guided denoising process and a trajectory stitching procedure to ensure feasibility. In contrast, Crowd-FM uses CFM to directly learn the transformation to Bernstein coefficients via ODE integration, enabling faster and more stable inference.
@@ III-B Flow Matching Objective
Following the independent coupling strategy employed in recent works like [5], a conditional probability path $p_{\tau}(\boldsymbol{\xi}|\boldsymbol{\xi_{0}},\boldsymbol{\xi_{1}})$ and a corresponding vector field $u_{\tau}(\boldsymbol{\xi}|\boldsymbol{\xi_{0}},\boldsymbol{\xi_{1}})$ is defined for any pair of samples $\boldsymbol{\xi_{0}}\sim q_{0}(\boldsymbol{\xi})$ and $\boldsymbol{\xi_{1}}\sim q_{1}(\boldsymbol{\xi}).$ This path is a simple linear interpolation in the high-dimensional space of control points:
@@ III-C Flow Matching Model Architecture
This subsection details the Conditional Flow Matching architecture (Fig. 2) adapted to learn the complex multi-modal distribution of Bernstein control points that define expert trajectories.
@@ III-C1 Multi-modal Conditioning Encoder
Point Cloud Encoder
For a point cloud input $X_{\mathrm{pcd}}\in\mathbb{R}^{N_{\mathrm{pts}}\times 2}$, where $N_{\mathrm{pts}}$ is the maximum number of point-cloud points, a feature extractor $f_{\mathrm{pcd}}$ composed of a series of 1D convolutions and ReLU activations maps each point to a high-dimensional feature vector. A global max-pooling operation is then applied across all points to yield a permutation-invariant global feature vector, $c_{\mathrm{pcd}}=\max_{i=1...N_{\mathrm{pts}}}f_{\mathrm{pcd}}(X_{\mathrm{pcd}})$. This vector is processed by a multi-layer perceptron (MLP), $g_{\mathrm{pcd}}$, to produce the final static embedding, $C_{\mathrm{static}}=g_{\mathrm{pcd}}(c_{\mathrm{pcd}})$.
Dynamic Obstacle Encoder
The Dynamic Obstacle information is compiled into a specialized tensor for the model to work with. For an input dynamic obstacle tensor $X_{\mathrm{dyn}}\in\mathbb{R}^{N_{\mathrm{obs}}\times 4}$, representing $N_{\mathrm{obs}}$, each with a 4D kinematic state(position and velocity), a dynamic obstacle encoder $f_{\mathrm{dyn}}$ processes them into a high-dimensional feature vector. An MLP converts obstacle features into high-dimensional embeddings, which are augmented with positional embeddings. Subsequently, a minimal Transformer Encoder with 4 heads and w layers employs self-attention to model interactions among the obstacles. Finally, max-pooling aggregates these features into a single vector, which a final MLP processes to produce the dynamic context embedding(D=512)(D=512), $C_{\mathrm{dynamic}}$.
Goal Encoder
The navigation goal is provided as a 2D heading vector, $X_{\mathrm{goal}}\in\mathbb{R}^{2}$. This vector is processed by a dedicated MLP, $g_{\mathrm{goal}}$, to produce a goal embedding: $C_{\mathrm{goal}}=g_{\mathrm{goal}}(\mathrm{X}_{\mathrm{goal}})$.
Multi-modal Fusion
The individual embeddings, $C_{\mathrm{static}}$, $C_{\mathrm{dyn}}$, and $C_{\mathrm{goal}}$, are treated as tokens for a fusion module. The tokens are concatenated to form an input sequence, which is then processed by $F_{\mathrm{fusion}}$, a Transformer Encoder with 8 heads and 3 layers. The Transformer Encoder uses self-attention to produce the final context, a sequence of fused tokens $C=F_{\mathrm{fusion}}(S)$. This sequence captures the complex interplay between the robot’s environment, other moving agents, and its objective. This context is directly used to condition the 1D U-Net as explained below.
@@ III-C2 Conditional U-Net
    The interpolated control point vector $\boldsymbol{\xi_{\tau}}\in\mathbb{R}^{2(n+1)}$.
@@
The U-Net architecture consists of a symmetric encoder-decoder structure with skip connections. The encoder progressively downsamples the 1D representation of the trajectory control points, and the decoder upsamples this representation, integrating information from the corresponding encoder layers via skip connections. The flow time $\tau$ is transformed into a sinusoidal time embedding and added to the intermediate features at each resolution level. The context $C$ is injected into the network within the U-Net’s residual blocks, allowing the model to condition the predicted flow on the specific environmental and goal context. The final output of the network is the predicted vector field $v_{\theta}(\boldsymbol{\xi_{\tau}},\tau,C)$, which has the same dimension as the input $\boldsymbol{\xi_{\tau}}$.
@@ III-D1 Guided Integration Process
It is important to note that $\mathbf{p}_{k}$ is a function of $\boldsymbol{\xi_{\tau}}$, and is implied that:
$\mathcal{L}_{\text{collision}}=f(\mathbf{p}_{k}),\quad\text{and}$
$\quad\mathbf{p}_{k}=g(\boldsymbol{\xi_{\tau}})\implies\mathcal{L}_{\text{collision}}=f(g(\boldsymbol{\xi_{\tau}}))$
@@ III-E2 Training Objective
The Scoring Function is trained using a cross-entropy loss, treating the problem as a K-class classification task, where the target labels are dynamically generated for each training instance. This objective trains the scorer to assign the highest logit to the candidate that best imitates the expert’s path. During inference, it is sufficient to select the trajectory index with the highest score.
@@
Therefore, while the absolute identity of any given index changes between training steps, the relative correspondence between the scores and the dynamically generated target is preserved, ensuring that the cross-entropy loss correctly penalizes the model for failing to assign the highest score to the geometrically optimal trajectory, as seen in Fig. 3.
@@ III-F Data Collection
TABLE I: Comparison of navigation success rates across different planners and environments with 35 dynamic agents. Note: CoHAN 2.0 success rates could be a little higher with more extensive parameter tuning.
Data collection was performed using the Barn environment to capture challenging static obstacle layouts and PEDSIM to model dense dynamic crowds. The dataset consists of around five hours of navigation data using both the Jackal and Turtlebot2 robots. Two sources of trajectories were included:
@@ IV VALIDATION AND BENCHMARKING
In this section, we detail our extensive evaluation of Crowd-FM. We show qualitative results, conduct quantitative benchmarking, and perform ablation studies in challenging simulated environments, comparing our method against three leading baselines: CrowdSurfer[7], DRL-VO[20], and CoHAN2.0[14]. We then validate the practical feasibility of our pipeline in real-world settings.
@@ IV-B Quantitative Analysis
Crowd-FM sets a new standard for performance in crowd navigation, outperforming all baselines in nearly every metric across standard PEDSIM[4] environments (Table I). While the performance of optimization-based planners like CoHAN 2.0 could potentially be improved through more exhaustive, environment-specific parameter tuning, Crowd-FM remains fundamentally superior in real-time inference efficiency, overall time-to-goal efficiency, and its inherent ability to prevent freezing behaviors in dense scenarios. This particularly improves upon CrowdSurfer, despite their shared use of generative priors. We stress-test both these pipelines in more challenging environments and share our results in Table II.
This improvement stems from two pillars: (1) our CFM model’s near-perfect learning of collision avoidance from data, and (2) exceptional computational speed (refer Sec. IV-B4). The following subsections are dedicated to ablation studies, especially focusing on the effectiveness of the vanilla CFM-generated trajectories, followed by the runtime analysis of our pipeline.
@@ IV-B1 Ablation Study without Optimizer Refinement
Here, we test Crowd-FM without the Optimizer refinement step by comparing the trajectories generated by the Vanilla-CFM with another baseline (DRL-VO) and see if the Conditional Flow Matching models the input data distribution well. As is evident from Table III, vanilla CFM alone is able to beat a leading RL-based baseline, DRL-VO, by a fair margin. This shows that CFM alone can model collision avoidance, especially in dense crowds. Qualitative examples of vanilla-CFM trajectories can be found in Fig. 5.
@@ IV-B2 Effect of Inference-time Cost Guidance
One of the two ways we employ inference-time refinement of CFM-generated trajectories is the Inference-time Cost Guidance. We test this property of Crowd-FM by disabling Optimizer refinement on the Cumberland and Freiburg environments in Pedsim with 25 agents for 10 trials. Table IV shows a clear improvement in Success rates for the collision-cost guided Flow objective compared to the un-guided one. Additionally, we present qualitative results for the same in Fig. 5.
Figure 5: Effect of Inference-time Cost Guidance: Left image shows the trajectories generated by Crowd-FM with Cost Guidance. The right image shows the trajectories generated by Crowd-FM without Cost Guidance. Trajectories generated using collision cost guidance are found to be more controlled near obstacles than the trajectories without it. (Tests done without Optimizer Refinement)
@@ IV-B3 Effect of the Learned Scoring Function
Lower HLP values indicate that the chosen trajectory is closer to the expert demonstration and thus more human-like. We compare trajectories selected by our learned Scoring Function against those chosen by predefined, hand-tuned cost functions that primarily optimize for collision avoidance and smoothness. These tests were done in an open-loop setting to take into account the expert human demonstrations. A custom wheelchair prototype was teleoperated in indoor environments amid dense crowds to create the test dataset. We evaluated HLPs for both methods on 10 such scenes and provided our results as a Grouped Bar Chart in Fig. 6.
@@
IV-B4 Runtime Analysis
To evaluate real-time feasibility, we benchmarked the Crowd-FM pipeline on a laptop with an NVIDIA RTX 3060 GPU. The entire planning loop, including data processing and trajectory generation, takes $\sim$75ms. A detailed breakdown shows that the CFM model generates the initial trajectory batch in $\sim$45ms, and the Learned Scoring Function selects the optimal candidate in $\sim$5ms. The remaining time is utilized by the Optimizer Refinement step ($\sim$25ms), which ensures dynamic feasibility. Performances are naturally expected to improve with better GPUs. Crowd-FM’s high computational efficiency maintains a consistent control rate suitable for dense crowd navigation.